# User Simulation for Evaluating Information Access Systems

### v1

@@ 6.2.2 Controlled Generation of Query Sequences
Next, we move from generating single queries to creating a set of queries with the same underlying search intent. The Controlled Query Generation method [Jordan:2006:JCDL] takes a set $R$ of documents as input, then generates a set of queries targeting these documents. Query generation is based on the identification of the most discriminating terms in the target document set using relative entropy. This way a query sequence $\langle q_{1},\dots,q_{n}\rangle$ can be created in a controlled way such that the first query ($q_{1}$) is theoretically the best and the last query $(q_{n})$ is theoretically the worst in identifying the source document set. Notice that unlike the previous probabilistic approaches, the query generation process here is deterministic in nature. Three types of queries are considered: (a) single term queries, (b) two term queries, and (c) queries of varying length. The specific steps are as follows:
  1. 1.
    Create a source set $R$ of documents, representing relevant results.
  2. 2.
    Construct a language model $P(t|\theta_{R})$ for the source document set by concatenating all documents $d\in R$ and treating them as a single piece of text.
  3. 3.
    Sort terms in the vocabulary as $t_{1},\dots,t_{m}$ based on their contribution to relative entropy:
    $score(t)=P(t|\theta_{R})\log\frac{P(t|\theta_{R})}{P(t|C)}~{},$ (6.3)
    where $P(t|C)$ is the collection (background) language model.
  4. 4.
    Use the most discriminating term $t_{1}$ to initiate the query generation process:
    1. (a)
      For single term queries, this term is used as the first query: $q_{1}=\{t_{1}\}$;
    2. (b)
      For two term queries, this term and the next largest contributor to relative entropy are used: $q_{1}=\{t_{1},t_{2}\}$;
    3. (c)
      For queries of varying length, the initial query is also the single most discriminating term: $q_{1}=\{t_{1}\}$.
  5. 5.
    Identify the next term $t_{i}$ that is the most discriminating and is not already in a query ($t_{i}\notin\{q_{1},\dots,q_{i-1}\}$).
  6. 6.
    Generate the next query $q_{i}$ using this term:
    1. (a)
      For single term queries, the next query is this term: $q_{i}=\{t_{i}\}$;
    2. (b)
      For two term queries, the next query is this term plus the most discriminating term: $q_{i}=\{t_{1},t_{i}\}$;
    3. (c)
      For queries of varying length, this term is concatenated to the previous query: $q_{i}=\{q_{i-1},t_{i}\}$.
  7. 7.
    Repeat Steps 5–6 until there are no more terms or the highest remaining term score is below some threshold ($score(t_{i})<\epsilon$).
@@ 6.2.3 Generating Query Reformulations
Baskaya:2012:SIGIR assume that a fixed set of terms $t_{1},\dots,t_{m}$ is available for each topic, from which queries may be constructed. Building on and extending the work of Keskustalo:2009:AIRS, they present five prototypical strategies, based on term level changes, that users may employ when modifying queries. These strategies are grounded in observed real life behaviour.
  • S1: an initial single-term query is followed by queries that repeatedly replace that term:
    ${}~{}~{}~{}~{}~{}~{}~{}q_{1}=\{t_{1}\}\rightarrow q_{2}=\{t_{2}\}\rightarrow q_{3}=\{t_{3}\}\rightarrow\dots$
  • S2: an initial two-term query is followed by queries repeatedly varying the second term:
    ${}~{}~{}~{}~{}~{}~{}~{}q_{1}=\{t_{1},t_{2}\}\rightarrow q_{2}=\{t_{1},t_{3}\}\rightarrow q_{3}=\{t_{1},t_{4}\}\rightarrow\dots$
  • S3: an initial three-term query is followed by queries repeatedly varying the third term:
    ${}~{}~{}~{}~{}~{}~{}~{}q_{1}=\{t_{1},t_{2},t_{3}\}\rightarrow q_{2}=\{t_{1},t_{2},t_{4}\}\rightarrow q_{3}=\{t_{1},t_{2},t_{5}\}\rightarrow\dots$
  • S4: an initial single-term query is followed by queries which extend the previous query with a new term:
    ${}~{}~{}~{}~{}~{}~{}~{}q_{1}=\{t_{1}\}\rightarrow q_{2}=\{t_{1},t_{2}\}\rightarrow q_{3}=\{t_{1},t_{2},t_{3}\}\rightarrow\dots$
  • S5: an initial two-term query is followed by queries which extend the previous query with a new term:
    ${}~{}~{}~{}~{}~{}~{}~{}q_{1}=\{t_{1},t_{2}\}\rightarrow q_{2}=\{t_{1},t_{2},t_{3}\}\rightarrow q_{3}=\{t_{1},t_{2},t_{3},t_{4}\}\rightarrow\dots$
@@
The above strategies assume that query terms are given in a predefined order and therefore the generation of query sequences is deterministic. To implement these approaches in practice, query terms may be selected and ordered by human annotators [Keskustalo:2009:AIRS], taken from the topic description by maintaining the original word order [Verberne:2015:ECIR], or sampled and ranked using a language model of the topic description [Maxwell:2016:CIKM].
@@
Carterette:2015:ICTIR generate queries dynamically, considering the history (queries and corresponding results) up to a given point $i$ in the session. It is assumed that topics $T=(s,Q)$ come with a textual description $s$ and a set of queries $Q$; such data is available, for example, in the TREC Session track [Carterette:2016:SIGIR]. One key component of the model is the conditioning of the query length $l$ on the topic, $P(l|T)$. This is motivated by the fact that some topics naturally lend themselves to longer queries. Another key element is that the language model from which query terms are sampled is continuously updated based on the results the user has seen for previous queries in the session. This is a simple way of modeling how users learn during the search process, which is disregarded in other query generation approaches. Instead of sampling queries from a traditional multinomial (unigram) language model, a two-stage process is used that yields more realistic queries. Queries are first generated by sampling from a series of binomial distributions, $P(t|T,l,i)$, then scored based on their discriminative power among topics, $P(q|T)$. Specifically, at each step $i$ of the session, the following steps are performed:
  1. 1.
    Generate $n$ candidate queries:
    • Sample query length $l$ according to $P(l|T)$.
    • Iterate over terms in $P(t|T,l,i)$ in order of decreasing probability:
      • Flip a coin to decide whether to add $t$ to the query.
      • Repeat until $l$ terms are sampled.
  2. 2.
    Sample one query from the set according to $P(q|T)$ to be returned as the simulated query reformulation $q_{i}$.
The basis of the query generation in Step 1) is a binomial model, $P(t|T,l,i)$, for term presence of queries of length $l$ for a topic. The maximum-likelihood estimate of this probability (based on the set of topic queries $Q$) could only recombine terms in previously seen queries, therefore it is smoothed with additional text data. For the first query of the session ($i=1$), additional terms come from the topic description $s$, while subsequent queries ($i>1$) consider terms from titles and snippets of documents seen by the user in previous results. Instead of sampling directly from this distribution, terms are iterated in descreasing order of probability and a binary decision (whether to add it to the query or not) is made for each term. As Carterette:2015:ICTIR explain, “this is meant to mitigate a problem with language models, that their probabilities tend to underestimate the importance of common terms while overestimating the importance of rare terms.”
Candidate queries in Step 2) are scored according to:
$P(q|T)=\prod_{t\in q}P(t|T)\propto\prod_{t\in q}P(T|t)P(t)~{},$ (6.4)
where the probability of a term given a topic, $P(t|T)$, measures the term’s ability to discriminate among topics. The estimate of this probability considers term appearances both in topic descriptions and in titles and snippets of documents retrieved for previous queries in the session, for the given topic and across topics. The prior probability of a term, $P(t)$, is treated as uniform.
@@ 6.2.5 User Knowledge and Cognitive State
Users with different backgrounds exhibit notable variations in their topical knowledge and prior experience with search engines. These factors influence the queries they would issue. Also, during the search session, users gain more knowledge about the topic as well as about how the search system works, both in general and with respect to their current information need, which influences future interactions. Currently, these factors either completely ignored or are modeled in a simple fashion, by representing users’ knowledge as term distributions [Maxwell:2016:CIKM, Camara:2022:ECIR].
@@ 6.3 Simulating Scanning Behaviour
Scanning behaviour is concerned with how the user processes the list of results presented to them in response to their search query. While the results are often displayed as a ranked list of snippets, more complex interfaces are also possible. From the perspective of modeling in the MDP framework, the user is in state $S_{S}$ when interacting with the presented search results. In this state, the actions that a user could potentially take $A_{S}$ depend on the interface features provided by the search system. For example, in a typical ranked list of snippets, a user can selectively examine some of the displayed snippets. The task of simulating scanning behaviour is to model the policy of a user for choosing which action to take based on the state $S_{S}$.
@@ 6.3.1 Models for Simulating Scanning Behaviour
Many models for simulating a user’s scanning behaviour have been proposed based on the sequential browsing hypothesis. Those models have often been developed as part of a clicking model for convenience of evaluation since it is hard to obtain the ground truth of a user’s scanning behaviour (which requires using devices such as eye-tracking), whereas the clicking actions of a user are easily observed and generally available in search log data. Since a clicking action can only happen after a user examines a snippet, when a result has been clicked, it implies that the user had examined that result (before clicking it). Under the sequential browsing assumption, one may also assume that any skipped results above a clicked result have also been examined by the user, but the user has just decided not to click on it (due to an unattractive snippet). However, whether a user has also examined a result ranked below the last clicked result would generally be unknown from the search log.
@@
Another interesting model is the user browsing model [Dupret:2008:SIGIR], which assumes sequential scanning behaviour. At each rank position, the user first decides whether to look at the snippet or not; if they find it attractive enough, then they click on it. Whether the result gets clicked or not, they will then resume the scan of the result list from the next rank position. Unlike the cascade model, this model does not assume that the user examines all result snippets up to a click. Instead, it models the event that user examines the snippet and, independently from it, whether they find the snippet attractive. Formally, the probability of result $i$ being examined is expressed as $P(E=1|R_{i},C_{1},...,C_{i-1})$, where $E$ is a binary random variable denoting the event of examination, $R_{i}$ is the result snippet at rank position $i$, and $C_{j}$ are a binary random variables indicating whether the result snippets at earlier rank positions got clicked. [^1] In [Dupret:2008:SIGIR] only the distance from the last clicked result is considered. The probability of result $R_{i}$ being attractive is expressed as $P(A=1|R_{i})$. Dupret:2008:SIGIR model both examination and attractiveness as Bernoulli variables. Neither $E$ nor $A$ are directly observed, hence the model parameters are estimated from clicks, which are modeled as a composition of these two events. The parameters of the examination and attractiveness models are estimated using the Expectation–Maximization (EM) algorithm. Their empirical results are consistent with the findings of the eye tracking studies of Joachims:2007:TOIS, namely: (1) the first two snippets receive substantially more user attention than later snippets, (2) the probability of examination or click decreases as the distance to the last click increases, and (3) users almost always look at the the snippet below a clicked snippet (referred to as “Skip Next” strategy in [Joachims:2007:TOIS]).
@@
In general, the parameters of a click model can be estimated based on search log data. When the click model involves only parameters representing observable events, the parameters can be easily estimated by normalizing the counts of the corresponding events of those parameters based on the analytical form of the Maximum Likelihood (ML) estimator. When a model involves a latent variable that not observable, which are clearly needed to model the scanning behaviour of a user, there is generally no analytical solution to the ML-estimation problem, and an iterative numerical algorithm such as the EM algorithm is often used to compute the ML estimate and estimate all the parameter values [Chuklin:2015:Book].
Note that the discussion above indicates the existence of complex dependencies between scanning behaviour and what kind of relevant information and how much relevant information the user has been able to gain in the process of interacting with the search results. This complexity coupled with the challenge in observing the actual scanning behaviour of a user makes the construction of an accurate model for simulating scanning behaviour quite challenging. As one solution, in the advanced models, a user’s scanning behaviour can be modeled via a latent variable parameter, which can then be estimated by fitting the model to the observed log data. While such an indirect way of modeling a user’s scanning behaviour could bypass the challenge to some extent, more user studies that target at directly studying a user’s scanning behaviour are clearly needed to make more substantial progress in simulating scanning behaviour.
@@ 6.3.2 Variations of Scanning Strategies
Klockner:2004:CHI observe that most users (52-65%) employ a depth-first strategy, where they click immediately on potentially relevant items as they encounter them. However, a non-negligible minority (11-15%) uses an breadth-first strategy and look through the entire result list, before clicking on any of the results. The remaining users (20-37%) apply a mixed partially breadth-first strategy, where they sometimes look ahead at the next few entries before selecting an item to click. Aula:2005:INTERACT identify two groups: economic (46%) and exhaustive (54%) evaluators. Economic evaluators tend to scan less than half of the results visible on the screen before their first action. Exhaustive evaluators tend to examine over half of the results visible on the screen, or even scroll down, before performing their first action. These groups resemble the depth-first and breadth-first strategies in [Klockner:2004:CHI]. Smucker:2011:HCIR distinguish between “fast and liberal” and “slow and neutral” searchers. Users in the former group are fast in examining snippets and liberally click on them, but rather spend more time evaluating the full documents. The latter group follows a reverse strategy: they take over twice as long to make a decision about which snippet to click to then speed up the evaluation of full documents. Dumais:2010:IIIX examine searcher behaviour when the SERP contains both organic results and advertisements. They identify three main clusters of users: those who explore the SERP broadly (exhaustive, 32%) and more narrowly (economic), with the latter category further broken down based on the attention given to advertisements (economic-results, 39% vs. economic-ads, 29%). Overall, two broad categories emerge from these studies, searchers who examine results exhaustively before making their first click and those who examine results more narrowly and click more liberally.
@@ 6.3.3 Complex Presentation Layouts
It is also important to remember that web search, while widely used, represents just one flavor of search. There are application domains where different result presentation formats are employed. A prime example is vertical search and recommender systems that focus on specific types of entities, such as products or videos, where a grid layout has emerged as the standard format for displaying search results. See Fig. 6.7 for examples of other result presentation layouts; in many of these, the top-down scanning assumption is misguided [Oosterhuis:2018:SIGIR]. Simulation of user behaviour enabled by complex SERPs with heterogeneous content and/or nonlinear page layouts requires the modeling of additional scanning, browsing, and clicking behaviours. How to formally model such complex browsing and interaction actions of a user is an important future direction. Given the complexity of such interfaces, model-based approaches might have inherent limitations compared to data-driven approaches [Azzopardi:2020:CHIIR], which goes back to the more fundamental question of model-based vs. data-driven simulation (cf. Section 1.4.1).

### v2

@@ 6.2.2 Controlled Generation of Query Sequences
Next, we move from generating single queries to creating a set of queries with the same underlying search intent. The Controlled Query Generation method [Jordan:2006:JCDL] takes a set $R$ of documents as input and generates a sequence of queries $\langle q_{1},\dots,q_{n}\rangle$ targeting these documents. This deterministic process identifies the most discriminating terms in the target documents using relative entropy, with the first query ($q_{1}$) theoretically being the best and the last query $(q_{n})$ the worst in identifying the source document set. A language model is generated from the documents in $R$, and terms are sorted based on their contribution to relative entropy. Query generation is initiated with the most discriminating term, and iteratively, the next most discriminating term not present in previous queries is identified to formulate the next query. The process terminates when no further terms remain or the highest term score falls below a predefined threshold. Three types of queries are considered: (a) single-term queries, (b) two-term queries, and (c) queries of varying length.
@@ 6.2.3 Generating Query Reformulations
Building on and extending the work of Keskustalo:2009:AIRS, Baskaya:2012:SIGIR present five prototypical query modification strategies based on observed user behaviour. It is assumed that a fixed set of terms ($t_{1},\dots,t_{m}$) is available for each topic, from which queries may be constructed. The strategies are based on term level changes that users may employ and include: (S1) repeated replacement of the initial term ($t_{1}\rightarrow t_{2}\rightarrow t_{3}\rightarrow\dots$), (S2) variation of the second term in a two-term query ($t_{1},t_{2}\rightarrow t_{1},t_{3}\rightarrow t_{1},t_{4}\rightarrow\dots$), (S3) variation of the third term in a three-term query ($t_{1},t_{2},t_{3}\rightarrow t_{1},t_{2},t_{4}\rightarrow t_{1},t_{2},t_{5}\rightarrow\dots$), (S4) extension of a single-term query with new terms ($t_{1}\rightarrow t_{1},t_{2}\rightarrow t_{1},t_{2},t_{3}\rightarrow\dots$), and (S5) extension of a two-term query with new terms ($t_{1},t_{2}\rightarrow t_{1},t_{2},t_{3}\rightarrow t_{1},t_{2},t_{3},t_{4}\rightarrow\dots$).
@@
These strategies assume that query terms are given in a predefined order and therefore the generation of query sequences is deterministic. To implement these approaches in practice, query terms may be selected and ordered by human annotators [Keskustalo:2009:AIRS], taken from the topic description by maintaining the original word order [Verberne:2015:ECIR], or sampled and ranked using a language model of the topic description [Maxwell:2016:CIKM].
@@
Carterette:2015:ICTIR generate queries dynamically, considering the history (queries and corresponding results) up to a given point $i$ in the session. It is assumed that topics $T=(s,Q)$ come with a textual description $s$ and a set of queries $Q$; such data is available, for example, in the TREC Session track [Carterette:2016:SIGIR]. One key component of the model is the conditioning of the query length $l$ on the topic, $P(l|T)$. This is motivated by the fact that some topics naturally lend themselves to longer queries. Another key element is that the language model from which query terms are sampled, $P(t|T,l,i)$, is continuously updated based on the results the user has seen for previous queries in the session. This is a simple way of modeling how users learn during the search process, which is disregarded in other query generation approaches. Instead of sampling queries from a traditional multinomial (unigram) language model, a two-stage process is used that yields more realistic queries. In the first step, multiple candidate queries are generated by sampling terms according to the binomial model $P(t|T,l,i)$. In the second step, one query is selected based on its discriminative power among topics, given by $P(q|T)$.
@@ 6.2.5 User Knowledge and Cognitive State
Users with different backgrounds exhibit notable variations in their topical knowledge and prior experience with search engines. These factors influence the queries they would issue. Also, during the search session, users gain more knowledge about the topic as well as about how the search system works, both in general and with respect to their current information need, which influences future interactions. Currently, these factors are either completely ignored or modeled in a simple fashion, by representing users’ knowledge as term distributions [Maxwell:2016:CIKM, Camara:2022:ECIR].
@@ 6.3 Simulating Scanning Behaviour
Scanning behaviour is concerned with how the user processes the list of results presented to them in response to their search query. While the results are often displayed as a ranked list of snippets, more complex interfaces are also possible. From the perspective of modeling in the MDP framework, the user is in state $S_{S}$ when interacting with the presented search results. In this state, the actions that a user could potentially take $A_{S}$ depend on the interface features provided by the search system. (To be consistent in notations, we tend to use variable $S$ to denote a state and variable $A$ to denote an action. The subscript $s$ here stands for “scanning.") For example, in a typical ranked list of snippets, a user can selectively examine some of the displayed snippets. The task of simulating scanning behaviour is to model the policy of a user for choosing which action to take based on the state $S_{S}$. Scanning behaviour is closely related to stopping behaviour as when viewed from decision perspective, modeling scanning behaviour includes modeling a user’s decision regarding whether to continue scanning or stop scanning. However, while the behaviour of continuing to scan a result list is well defined, the stopping behaviour is more complicated as there may be multiple reasons for stopping, including, e.g., when a user’s information need is already satisfied, when a user has given up search, or when a user has decided to reformulate the query. For this reason, we have decided to discuss the stopping behaviour in a separate section later (see Section 6.3).
@@ 6.3.1 Models for Simulating Scanning Behaviour
Many models for simulating a user’s scanning behaviour have been proposed based on the sequential browsing hypothesis. The C/W/L framework [Moffat:2013:CIKM, Moffat:2017:TOIS] and its extension, the C/W/L/A framework [Moffat:2022:SIGIR] are particularly interesting because a variable $C(i)$ is introduced to explicitly model the probability of continuing to scan after examining an item at position $i$, referred to as a browsing model in [Moffat:2022:SIGIR]. This enables instantiating the evaluation framework with any specific browsing model, which can be regarded as the same as a scanning model. Similarly, in the work [Carterette:2011:SIGIR], a general conceptual framework also explicitly identified browsing model as a separate component (enabling different instantiations), along with two other components, i.e., model of document utility, and utility accumulation model.
Specific scanning/browsing models have often been developed as part of a clicking model for convenience of evaluation since it is hard to obtain the ground truth of a user’s scanning behaviour (which requires using devices such as eye-tracking), whereas the clicking actions of a user are easily observed and generally available in search log data. Since a clicking action can only happen after a user examines a snippet, when a result has been clicked, it implies that the user had examined that result (before clicking it). Under the sequential browsing assumption, one may also assume that any skipped results above a clicked result have also been examined by the user, but the user has just decided not to click on it (due to an unattractive snippet). However, whether a user has also examined a result ranked below the last clicked result would generally be unknown from the search log.
@@
Another interesting model is the user browsing model [Dupret:2008:SIGIR], which assumes sequential scanning behaviour. At each rank position, the user first decides whether to look at the snippet or not; if they find it attractive enough, then they click on it. Whether the result gets clicked or not, they will then resume the scan of the result list from the next rank position. Unlike the cascade model, this model does not assume that the user examines all result snippets up to a click. Instead, it models the event that user examines the snippet and, independently from it, whether they find the snippet attractive. Formally, the probability of result at rank position $i$ being examined is expressed as $P(E=1|R_{i},C_{1},...,C_{i-1})$, where $E$ is a binary random variable denoting the event of examination, $R_{i}$ is the result snippet at rank position $i$, and $C_{j}$ are a binary random variables indicating whether the result snippets at earlier rank positions got clicked. [^3] In [Dupret:2008:SIGIR] only the distance from the last clicked result is considered. The probability of result $R_{i}$ being attractive is expressed as $P(A=1|R_{i})$. Dupret:2008:SIGIR model both examination and attractiveness as Bernoulli variables. Neither $E$ nor $A$ are directly observed, hence the model parameters are estimated from clicks, which are modeled as a composition of these two events. The parameters of the examination and attractiveness models are estimated using the Expectation–Maximization (EM) algorithm. Their empirical results are consistent with the findings of the eye tracking studies of Joachims:2007:TOIS, namely: (1) the first two snippets receive substantially more user attention than later snippets, (2) the probability of examination or click decreases as the distance to the last click increases, and (3) users almost always look at the the snippet below a clicked snippet (referred to as “Skip Next” strategy in [Joachims:2007:TOIS]).
@@
Note that the discussion above indicates the existence of complex dependencies between scanning behaviour and what kind of relevant information and how much relevant information the user has been able to gain in the process of interacting with the search results. This complexity coupled with the challenge in observing the actual scanning behaviour of a user makes the construction of an accurate model for simulating scanning behaviour quite challenging. Since it is impossible to observe the latent decision/preference behaviour of a user while scanning through the results, we have not yet seen results to examine to what extent a scanning behaviour model can capture a real user’s scanning behaviour. As one solution, in the advanced models, a user’s scanning behaviour can be modeled via a latent variable parameter, which can then be estimated by fitting the model to the observed log data. While such an indirect way of modeling a user’s scanning behaviour could bypass the challenge to some extent, more user studies that target at directly studying a user’s scanning behaviour are clearly needed to make more substantial progress in simulating scanning behaviour.
@@ 6.3.2 Variations of Scanning Strategies
Klockner:2004:CHI observe that most users (52-65%) employ a depth-first strategy, where they click immediately on potentially relevant items as they encounter them. However, a non-negligible minority (11-15%) uses an breadth-first strategy and look through the entire result list, before clicking on any of the results. The remaining users (20-37%) apply a mixed partially breadth-first strategy, where they sometimes look ahead at the next few entries before selecting an item to click. Aula:2005:INTERACT identify two groups: economic (46%) and exhaustive (54%) evaluators. Economic evaluators tend to scan less than half of the results visible on the screen before their first action. Exhaustive evaluators tend to examine over half of the results visible on the screen, or even scroll down, before performing their first action. These groups resemble the depth-first and breadth-first strategies in [Klockner:2004:CHI]. Smucker:2011:HCIR distinguish between “fast and liberal” and “slow and neutral” searchers. Users in the former group are fast in examining snippets and liberally click on them, but rather spend more time evaluating the full documents. The latter group follows a reverse strategy: they take over twice as long to make a decision about which snippet to click to then speed up the evaluation of full documents. Dumais:2010:IIIX examine searcher behaviour when the SERP contains both organic results and advertisements. They identify three main clusters of users: those who explore the SERP broadly (exhaustive, 32%) and more narrowly (economic), with the latter category further broken down based on the attention given to advertisements (economic-results, 39% vs. economic-ads, 29%). Overall, two broad categories emerge from these studies: searchers who examine results exhaustively before making their first click and those who examine results more narrowly and click more liberally.
@@ 6.3.3 Complex Presentation Layouts
It is also important to remember that web search, while widely used, represents just one flavor of search. There are application domains where different result presentation formats are employed. A prime example is vertical search and recommender systems that focus on specific types of entities, such as products or videos. Vertical search often employs a grid layout to display search results, while many recommender systems employ a carousel-based presentation format [Bendada:2020:RecSys, Rahdari:2022:HT]. See Fig. 6.8 for examples of various result presentation layouts; in many of these, the top-down scanning assumption is misguided [Oosterhuis:2018:SIGIR]. Simulation of user behaviour enabled by complex SERPs with heterogeneous content and/or nonlinear page layouts requires the modeling of additional scanning, browsing, and clicking behaviours. How to formally model such complex browsing and interaction actions of a user is an important future direction. There is ongoing work in this area, for example, on user interaction models for carousel-based layouts for recommender systems [Felicioni:2021:UMAP, Rahdari:2024:TORS]. However, given the increasing complexity of result presentation interfaces, model-based approaches might have inherent limitations compared to data-driven approaches [Azzopardi:2020:CHIIR], which goes back to the more fundamental question of model-based vs. data-driven simulation (cf. Section 1.4.3).